You are given the data below for 2008 for the imaginary country of Amagre, whose currency is the G.


Consumption 350 billion G

Transfer payments 100 billion G

Investment 100 billion G

Government purchases 200 billion G

Exports 50 billion G

Imports 150 billion G

Bond purchases 200 billion G

Earnings on foreign investments 75 billion G

Foreign earnings on Amagre investment 25 billion G

Compute net foreign investment.

Compute net exports.

Compute GDP.

Compute GNP.


1
Expert's answer
2021-10-08T17:12:22-0400

Net foreign investment is:

NFI = 75 - 25 = 50 billion G.

Net exports is:

NX = X - M = 50 - 150 = -100 billion G.

GDP = C + I + G + NX = 350 + 100 + 200 - 100 = 550 billion G.

GNP = GDP - NFI = 550 - 50 = 500 billion G.